A vendor mixed 3 litres of water into 17 litres of milk what is the percentage of milk in the mixture

Grade:8

2 Answers

RIZWAN IQBAL
2 years ago
The percentage of the milk can be calculated be just multiplying 17 with 5 because the total mixture quantity is 20,so 20x 5 will make it 100,hence making it very easy to make out what is the percentage,so 17x 5 =85,this means that the percentage of milk In the mixture is 85%...hope this solved your doubt.
Debarati
2 years ago
Amount of milk= 17 litres
Amount of water= 3 litres
Amount of total mixture= 17+3= 20 litres
Percentage of milk in the mixture= (17/20)*100= 0.85*100= 85%